# Minimum Strict Consistent Subset in Paths, Spiders, Combs and Trees

### v2

@@ 3.2 NP-hardness of the MSCS in General Graphs
Our reduction is from dominating set problem on a graph $G=(V,E)$. From $G$, we construct another graph $G^{\prime}=G_{1}\cup G_{2}\cup\{v_{spc}\}$ where $G_{1}$ and $G_{2}$ are the copies of $G$ and $v_{spc}$ is a singleton node. Each edge of $G_{1}$ and $G_{2}$ has weight $1$. Every vertices of $G_{1}$ are directly connected by edges with all the vertices of $G_{2}$ with weights of edge $3-4\epsilon$ where $0<\epsilon<\frac{1}{2}$. Additionally, all vertices of $G_{2}$ are directly connected by edges of weight $2\epsilon$ to $v_{spc}$. All the vertices of $G_{1}$ are colored red and all the vertices of $G_{2}$ are colored blue and $v_{spc}$ is also colored blue. This completes the construction.
@@
Suppose that $G$ has a dominating set $S$ of size $k$. We will show that $G^{\prime}$ has a strict consistent subset of size $k+1$. Suppose that $S^{\prime}\leftarrow\emptyset$. For each vertex of $S$, we put the corresponding vertex in $G_{1}$ in $S^{\prime}$. Also, we put $v_{spc}$ in $S^{\prime}$. Therefore, $\lvert S^{\prime}\rvert=k+1$. We claim that $S^{\prime}$ is a strict consistent subset of $G^{\prime}$. For a red color vertex $v\in G_{1}$, if $v\notin S^{\prime}$, then $\mathrm{d}(v,v_{spc})=3-4\epsilon+2\epsilon>1$, but $v$ has a neighbor vertex in $G_{1}$ which is in $S^{\prime}$ of length $1$. Therefore, all the red-colored vertices are strictly covered. Also, for a blue-colored vertex $u\in G_{2}$, $\mathrm{d}(u,G_{1})=3-4\epsilon>2\epsilon=\mathrm{d}(u,v_{spc})$. Therefore, all the red-colored vertices are strictly covered by $v_{spc}$.
Again, suppose that $S^{\prime}$ is a strict consistent subset of $G^{\prime}$ such that $\lvert S^{\prime}\rvert=k+1$. Therefore $v_{spc}$ must included in $S^{\prime}$; otherwise, size of $S^{\prime}$ must be greater than $k+1$. Therefore, all the vertices of $G_{2}$ are strictly covered. Hence we have $k$ vertices from $G_{1}$ in $S^{\prime}$. We take corresponding vertices of $G_{1}$ in $S^{\prime}$ into a set $S$. We claim that $S$ is the dominating set of $G$. If not then there exists a vertex $v\in S$ such that there is no neighbor of $v$ in $S$ and $v\notin S$. Therefore the corresponding vertex of $v$ in $G_{1}$ can not be covered strictly by $S^{\prime}$; a contradiction. ∎
@@
Recall the block tree $\mathcal{B}(\mathrm{T})$; see Figure 2. Let $S$ be the empty set. For every edge $e\in\mathcal{B}(\mathrm{T})$ our algorithm adds to $S$ the two vertices of $\mathrm{T}$ that correspond to the endpoints of $e$. It is easily seen that $S$ is a feasible solution for the MSCS problem. We claim that $S$ is a 2-approximate solution. Let $b$ denote the number of blocks in $\mathrm{T}$, which is the same as the number of vertices of $\mathcal{B}(\mathrm{T})$. Since $\mathcal{B}(\mathrm{T})$ is a tree, it has $b-1$ edges. Therefore $\lvert S\rvert=2b-2$. On the other hand the size of any optimal solution $OPT$ is at least $b$ because it must contain at least one vertex from each block of $\mathrm{T}$ (by Lemma 3.1), and hence $\lvert OPT\rvert\geq b$. Therefore
@@
We note that this elementary analysis is the best possible for this algorithm, for example if $\mathrm{T}$ is a path with blocks of size three, then our algorithm picks two vertices for each block (except for the two leaf blocks) while the optimal solution picks one vertex (the middle vertex) from each block.
@@ 4 Algorithm for Finding MSCS in Trees
The Lemma 3.1 clearly demonstrates that at least one vertex from each block must appear in any strict consistent subset which is the restriction of the paper [14]. However, the definitions of minimum consistent spanning subset and minimum strict consistent subset are something distinct, but the Lemma [14] is applicable for these two variants of MCS. Therefore, we use a very similar argument from the paper [14] as follows for our MSCS problem.
Lemma 3.2 suggests a more constrained version of the MSCS problem, in the sense that we can fix a leaf block B and enforce exactly one vertex of B to be in the solution. As we do not know in advance which vertex of B is in the optimal solution, we try all of them and report the best answer.
Our algorithm employs a nontrivial dynamic programming approach. First we introduce the subproblems that will be generated throughout the algorithm and then we will show how to solve the subproblems recursively.
4.1 Defining Subproblems For Solving MSCS in Trees
We denote each subproblem by $\mathrm{T}(x,y)$ where $x$ and $y$ are two given vertices of $\mathrm{T}$. Consider $\mathrm{path}(x,y)$ between $x$ and $y$ in $\mathrm{T}$ and let $u\in\mathrm{N}(y)$ in $\mathrm{path}(x,y)$ (it might be the case that $x$ = $u$). By removing the edge $(u,y)$ from $\mathrm{T}$ we obtain two subtrees. Let $\mathrm{T}_{y}$ be the subtree containing $y$ (see Figure 4(A)). Let $\mathrm{T}^{\prime}$ be the union of $\mathrm{path}(x,y)$ and $\mathrm{T}_{y}$ (see Figure 4(B)). We define $\mathrm{T}(x,y)$ to be the MSCS problem on $\mathrm{T}^{\prime}$ with the following constraints:
  • $x$ must be in the solution, and
  • all the vertices from $x$ to $u$ on $\mathrm{path}(x,y)$ must be covered by $x$.
These constraints imply that the vertices from $x$ to $u$ should have the same color.
@@ 4.1 Defining Subproblems For Solving MSCS in Trees
4.2 Solving Subproblems For Solving MSCS in Trees
@@
If $\mathrm{T}^{\prime}$ is monochromatic, then (by Observation 3.0.1) we return $x$ as the solution. In this case, $S(x,y)=1$. Assume that $\mathrm{T}^{\prime}$ is not monochromatic. We root $\mathrm{T}^{\prime}$ at $x$.
@@
If $\mathrm{T}(x,y)$ has a solution, then any solution of $\mathrm{T}(x,y)$ contains a vertex $z$ in the same block as $x$ or in a neighboring block of $x$ such that all vertices on $\mathrm{path}(x,z)$ are covered by $x$ or by $z$.
@@
As $\mathrm{T}^{\prime}$ is multicolored, any solution of $\mathrm{T}(x,y)$ should contain at least two vertices. In particular it should contain at least one vertex from each neighboring block of $x$. In any solution of $\mathrm{T}(x,y)$ a vertex that is closest to $x$ satisfies the statement of the lemma for $z$. ∎
Let $z$ be any vertex of $\mathrm{T}^{\prime}$ that satisfies the constraints of Lemma 4.1 (see Figure 5) (It might be the case that $z=y$. Also $z$ could be in $x$’s block or in $x$’s neighboring block.) If such a vertex $z$ does not exist then $\mathrm{T}(x,y)$ has no solution and thus we set $S(x,y)=+\infty$.
@@
By Lemma 4.1 all vertices of $\mathrm{path}(x,z)$ are covered by $x$ or $z$. Since $u$ must be covered by $x$ (as imposed by the definition of $\mathrm{T}(x,y))$, we must have $\mathrm{d}(u,x)<\mathrm{d}(u,z)$, and thus $\mathrm{d}(z,x)>2\cdot\mathrm{d}(u,x)$. Moreover as described before $\mathrm{path}(x,z)$ has $2k$ vertices (including a and z). Therefore, the first $k$ vertices must be covered by $x$ and the second $k$ vertices by $z$.
@@
Now we are in a problem setting where both $x$ and $z$ must be in the solution, and all vertices on $\mathrm{path}(x,z)$ must be covered by $x$ and $z$. We denote this more constrained version of $\mathrm{T}(x,y)$ by problem $\mathrm{T}(x,y,z)$ (we do not call this a subproblem for a reason to be determined later). In other words $\mathrm{T}(x,y,z)$ is the MSCS problem on $\mathrm{T}^{\prime}$ with the following constraints:
  • $z$ is in the same block as $x$ or in a neighboring block of $x$,
  • $x$ and $z$ must be in the solution,
  • $\mathrm{d}(z,x)>2\cdot\mathrm{d}(u,x)$, and
@@
  • $\mathrm{path}(x,z)$ has $2k$ where the first k vertices must have the same color as $x$ and the second $k$ vertices must have the same color as $z$.
@@
We refer to any vertex $z$ that satisfies the above constraints as a valid pair for $x$. Now we show how to solve $\mathrm{T}(x,y,z)$. We denote the size of the solution for $\mathrm{T}(x,y,z)$ by $S(x,y,z)$. Let $X$ be the set of all the vertices on $\mathrm{path}(y,z)$ that are closer to $x$ than to $z$ as in Figure 5. Let $Z$ be the set of all the vertices on $\mathrm{path}(y,z)$ that are closer to $z$ than to $x$.
@@
To solve $\mathrm{T}(x,y,z)$ we define two sets $\mathcal{X}$ and $\mathcal{Z}$ as follows. For each vertex $v$ in $X$, we add to $\mathcal{X}$ all children of $v$ that are not on the path $\mathrm{path}(y,z)$. For each vertex $v$ in $Z$, we add to $\mathcal{Z}$ all children of $v$ that are not on the path $\mathrm{path}(y,z)$. Then, the solution of $\mathrm{T}(x,y,z)$ is obtained by taking the union of $\{x,z\}$ with the solutions of $\mathrm{T}(x,v^{\prime})$ for all $v^{\prime}\in\mathcal{X}$ and the solutions of $\mathrm{T}(z,v^{\prime})$ for all $v^{\prime}\in\mathcal{Z}$. It might happen that some choices of $z$ does not lead to a valid solution, but then a subsequent subproblem, e.g. $\mathrm{T}(z,v^{\prime})$ will be infeasible and $S(z,v^{\prime})=+\infty$.
@@
One might wonder that for two vertices $v^{\prime}_{1},v^{\prime}_{2}\in\mathcal{Z}$ the solutions of $\mathrm{T}(z,v^{\prime}_{1})$ and $\mathrm{T}(z,v^{\prime}_{2})$ may affect each other. However, this cannot happen because by the definition of $\mathrm{T}(.,.)$ all vertices on paths from $z$ to the parents of $v^{\prime}_{1}$ and $v^{\prime}_{2}$ must be covered by $z$, and thus any vertex in the solutions of $\mathrm{T}(z,v^{\prime}_{1})$ and $\mathrm{T}(z,v^{\prime}_{2})$ lies in the same level as $z$ or in a lower level (i.e. higher depth) in $\mathrm{T}^{\prime}$. The same argument applies to vertices in $\mathcal{X}$. Notice that no vertex has an equal distance to multiple vertices of different colors in the final solution because of the definition of MSCS. Therefore,
@@
Now we turn our attention back to the original problem $\mathrm{T}(x,y)$. If we knew $z$ then the solution of $T(x,y,z)$ would also be a solution for $T(x,y)$. But we do not know $z$ in advance. Therefore we consider all valid pairs $z$ for $x$ and take one that gives the smallest solution. Let $P$ denote the set of all valid pairs for $x$. Then
@@
If $P$ is the empty set, i.e. $x$ has no valid pairs, then we set $S(x,y)=+\infty$. This is the end of our solution for the subproblem $\mathrm{T}(x,y)$; the problem $T(x,y,z)$ was introduced to simplify the description of our recursive solution.
@@ 4.3 Solving the Original Problem in Trees
In this section we show how to solve the original MSCS problem on the tree $\mathrm{T}$. Let $S$ denotes an optimal solution for this problem. If $\mathrm{T}$ is monochromatic then we return an arbitrary vertex as a solution. Assume that $\mathrm{T}$ has more than one colors. Let $L$ be a leaf block of $tree$, and let $\mathrm{N}$ be the (only) neighboring block of $L$. Let $y$ be the vertex of $\mathrm{N}$ adjacent to a vertex of $L$. By Lemma 3.2, any optimal solution has exactly one vertex, say $x$, from L. Thus, $x$ is the closest vertex of $S$ to all vertices in $L$. Hence the original problem is an instance of the $\mathrm{T}(x,y)$ problem that was introduced earlier. Since we do not know $x$, we try all vertices of $L$. Therefore,

### v3

@@ 3.2 NP-hardness of the MSCS in General Graphs
We reduce our problem from the dominating set problem on graph $G$. We create a graph $G^{\prime}=G_{1}\cup G_{2}\cup\{v_{sng}\}$ from $G$, where $G_{1}$ and $G_{2}$ are copies of $G$ and $v_{sng}$ is a singleton vertex. Each edge of $G_{1}$ and $G_{2}$ has weight $1$. All vertices of $G_{1}$ are adjacent to all vertices of $G_{2}$ by edges. The weight of the edges between $G_{1}$ and $G_{2}$ is $3-4\epsilon$, where $0<\epsilon<\frac{1}{2}$. Additionally, all vertices of $G_{2}$ are directly connected by edges of weight $2\epsilon$ to $v_{sng}$ (Figure 3(B)). Each vertex in $G_{1}$ is red color, each vertex in $G_{2}$ is green color, and $v_{sng}$ is also green color. This completes the construction of our reduction.
@@
Suppose that $G$ has a dominating set $S$ of size $k$. We will show that $G^{\prime}$ has a strict consistent subset of size $k+1$. Suppose that $S^{\prime}\leftarrow\emptyset$. For each vertex of $S$, we put the corresponding vertex in $G_{1}$ in $S^{\prime}$. Also, we put $v_{sng}$ in $S^{\prime}$. Therefore, $\lvert S^{\prime}\rvert=k+1$. We claim that $S^{\prime}$ is a strict consistent subset of $G^{\prime}$. For a red color vertex $v\in G_{1}$, if $v\notin S^{\prime}$, then $\mathrm{d}(v,v_{sng})=3-4\epsilon+2\epsilon>1$, but $v$ has a neighbor vertex in $G_{1}$ which is in $S^{\prime}$ of length $1$. Therefore, all the red-colored vertices are strictly covered. Also, for a green-colored vertex $u\in G_{2}$, $\mathrm{d}(u,G_{1})=3-4\epsilon>2\epsilon=\mathrm{d}(u,v_{sng})$. Therefore, all the red-colored vertices are strictly covered by $v_{sng}$.
Again, suppose that $S^{\prime}$ is a strict consistent subset of $G^{\prime}$ such that $\lvert S^{\prime}\rvert=k+1$. Therefore $v_{sng}$ must included in $S^{\prime}$; otherwise, size of $S^{\prime}$ must be greater than $k+1$. Therefore, all the vertices of $G_{2}$ are strictly covered. Hence we have $k$ vertices from $G_{1}$ in $S^{\prime}$. We take corresponding vertices of $G_{1}$ in $S^{\prime}$ into a set $S$. We claim that $S$ is the dominating set of $G$. If not then there exists a vertex $v\in S$ such that there is no neighbor of $v$ in $S$ and $v\notin S$. Therefore the corresponding vertex of $v$ in $G_{1}$ can not be covered strictly by $S^{\prime}$; a contradiction. ∎
@@
Consider the block tree $\mathcal{B}(\mathrm{T})$, as shown in Figure 1. Let $S$ be an empty set. Our technique adds the two vertices of $\mathrm{T}$ that correspond to the endpoints of $e$ to $S$ for every edge $e\in\mathcal{B}(\mathrm{T})$. Clearly, $S$ is a feasible solution for the MSCS problem. We assert $S$ is a 2-approximate solution. Let $b$ represent the number of blocks in $\mathrm{T}$, which is the same as the number of vertices in $\mathcal{B}(\mathrm{T})$. As $\mathcal{B}(\mathrm{T})$ is a tree, it has $b-1$ edges. Therefore, $\lvert S\rvert=2b-2$. However, any optimal solution $OPT$ must contain at least one vertex from each block of $\mathrm{T}$ (Lemma 3.1). So, $\lvert OPT\rvert\geq b$. Therefore
@@
This elementary analysis is the best for this algorithm. For example, if $\mathrm{T}$ is a path with $b$ blocks, each with three vertices, our algorithm selects two vertices (except the two leaf blocks), and the optimal solution selects the middle vertex from each block. Each leaf block has exactly one vertex in $S$. Therefore, $\lvert OPT\rvert=b$ and $\lvert S\rvert=2b-2$, proving our assertion.
@@ 4 Algorithm for Finding MSCS in Trees
According to Lemma 3.2, the MSCS problem can be more easier by taking a leaf block $B$ and requiring only one vertex to be in the solution. We try all the vertices of $B$ and report the best one.
We use nontrivial dynamic programming in our algorithm. We will first present the subproblems, and then we show how to solve subproblems recursively.
4.1 Defining Subproblems in Trees
Each subproblem is represented as $\mathrm{T}(x,y)$, where $x$ and $y$ are two vertices of $\mathrm{T}$. Consider $\mathrm{path}(x,y)$ between $x$ and $y$ in $\mathrm{T}$. Let $u\in\mathrm{N}(y)$ in $\mathrm{path}(x,y)$ (possibly $x=u$). Removing the edge $(u,y)$ from $\mathrm{T}$ yields two subtrees. Refer to Figure 4(A) for the subtree containing $y$ known as $\mathrm{T}_{y}$. Let $\mathrm{T}^{\prime}$ be the union of $\mathrm{path}(x,y)$ and $\mathrm{T}_{y}$ (see to Figure 4(B We define $\mathrm{T}(x,y)$ as the MSCS problem on $\mathrm{T}^{\prime}$ with the following constraints:
  • the solution must include the vertex $x$, and
  • $x$ must cover all vertices on $\mathrm{path}(x,y)$, from $x$ to $u$.
It is clear from the above constraints that $x$ to $u$ should have the same color according to the definition of MSCS.
@@ 4.1 Defining Subproblems in Trees
4.2 Solving the Subproblems in Trees
@@
Observation 3.0.1 states that if $\mathrm{T}^{\prime}$ is monochromatic, then we return $x$ as a solution. In this situation, $S(x,y)=1$. Assume $\mathrm{T}^{\prime}$ isn’t monochromatic. Now, $x$ is the root of $\mathrm{T}^{\prime}$.
@@
Suppose $\mathrm{T}(x,y)$ has a solution. Any solution of $\mathrm{T}(x,y)$ has a vertex $z$ in the same or neighboring block of $x$ such that $x$ or $z$ covers all vertices on $\mathrm{path}(x,z)$.
@@
Any solution to $\mathrm{T}(x,y)$ must have at least two vertices, as $\mathrm{T}^{\prime}$ is multicolored. In particular, it should have a vertex from each neighboring block of $x$. Therefore, for such a vertex $z$, the lemma is satisfied by the vertex closest to $x$ in any solution of $\mathrm{T}(x,y)$. ∎
Let $z$ be any vertex of $\mathrm{T}^{\prime}$ that satisfy the Lemma 4.1 (see Figure 5). It also case that $z=y$ may happen. Also, $z$ could be in $x$’s block or its neighbor. If a vertex $z$ does not exist, $\mathrm{T}(x,y)$ has no solution, hence $S(x,y)=+\infty$.
@@
Lemma 4.1 states that $x$ or $z$ covers all vertices of $\mathrm{path}(x,z)$. The definition of $\mathrm{T}(x,y)$ requires that $u$ be covered by $x$, thus $\mathrm{d}(u,x)<\mathrm{d}(u,z)$ and $\mathrm{d}(z,x)>2\cdot\mathrm{d}(u,x)$. Furthermore, if $x$ and $z$ have different colors, $\mathrm{path}(x,z)$ has an even number of vertices (see Figure 6(A)); otherwise, it does not satisfy the definition of MSCS. Assume $\mathrm{path}(x,z)$ has $2k$ vertices. The first $k$ vertices must be covered by $x$, and the second $k$ by $z$. If $x$ and $z$ have different colors, then it does not matter whether length of $\mathrm{path}(x,z)$ is even or odd (see Figure 6(B)).
Figure 6: (A) The tree $\mathrm{T}$ , and (B) the tree $\mathrm{T}^{\prime}$.
@@
Now we see that both $x$ and $z$ must be present in the solution, and they must cover all vertices on $\mathrm{path}(x,z)$. We represent this more constrained version of $\mathrm{T}(x,y)$ as $\mathrm{T}(x,y,z)$ (we do not call this a subproblem for a reason to be determined later). In other words, $\mathrm{T}(x,y,z)$ represents the MSCS problem on $\mathrm{T}^{\prime}$ with the following constraints:
  • $z$ is in the same block as $x$ or in a neighboring block of $x$.
  • The solution must include both $x$ and $z$.
  • $\mathrm{d}(z,x)>2\cdot\mathrm{d}(u,x)$.
@@
  • $\mathrm{path}(x,z)$ has $2k$ where the first $k$ vertices must have the same color as $x$ and the second $k$ vertices must have the same color as $z$ when $C(x)\neq C(z)$.
  • If $C(x)=C(z)$, then the length of the path $\mathrm{path}(x,z)$ can be even or odd. If length of the path $\mathrm{path}(x,z)$ is odd then the middle vertex is covered by anyone of $x$ and $z$ (see Figure 6(B)).
@@
A valid pair for $x$ is defined as any vertex $z$ that satisfies the above constraints. Now we demonstrate how to solve $\mathrm{T}(x,y,z)$. The size of the solution for $\mathrm{T}(x,y,z)$ is denoted as $S(x,y,z)$. Let $X$ be the set of all vertices on $\mathrm{path}(y,z)$ that are closer to $x$ than to $z$, as shown in Figure 5. Let $Z$ denote the set of all vertices on $\mathrm{path}(y,z)$ that are closer to $z$ than to $x$.
@@
To solve $\mathrm{T}(x,y,z)$, we define two sets $\mathcal{X}$ and $\mathcal{Z}$ as shown below. For each vertex $v$ in $X$, we add to $\mathcal{X}$ those children of $v$ who are not on the path $\mathrm{path}(y,z)$. For each vertex $v$ in $Z$, we add to $\mathcal{Z}$ all children of $v$ who are not on the path $\mathrm{path}(y,z)$. The solution of $\mathrm{T}(x,y,z)$ is found by taking the union of $\{x,z\}$ with the solutions of $\mathrm{T}(x,v^{\prime})$ for all $v^{\prime}\in\mathcal{X}$ and the solutions of $\mathrm{T}(z,v^{\prime})$ for all $v^{\prime}\in\mathcal{Z}$. It might be a case that not all choices of $z$ may result in a valid solution. In such a case, $\mathrm{T}(z,v^{\prime})$ will be infeasible, and $S(z,v^{\prime})=+\infty$.
@@
For two vertices $v^{\prime}_{1},v^{\prime}_{2}\in\mathcal{Z}$, the solutions of $\mathrm{T}(z,v^{\prime}_{1})$ and $\mathrm{T}(z,v^{\prime}_{2})$ may affect each other. However, this cannot happen because by the definition of $\mathrm{T}(.,.)$ all vertices on paths from $z$ to the parents of $v^{\prime}_{1}$ and $v^{\prime}_{2}$ must be covered by $z$, and thus any vertex in the solutions of $\mathrm{T}(z,v^{\prime}_{1})$ and $\mathrm{T}(z,v^{\prime}_{2})$ lies in the same level as $z$ or in a lower level (i.e. higher depth) in $\mathrm{T}^{\prime}$. The same argument applies to vertices in $\mathcal{X}$. In the final solution, no vertex has an equal distance to multiple vertices of different colors due to the definition of MSCS. Therefore,
@@
Now we return to the original problem, $\mathrm{T}(x,y)$. If we knew $z$, we can solve $T(x,y,z)$ as well as $T(x,y)$. However, we do not know $z$ in advance. As a result, we consider all valid $z$ combinations for $x$ and select the one that gives us the smallest solution. Let $P$ be the collection of all valid pairs for $x$. Then
@@
If $P$ is the empty set, which means $x$ has no valid pairs, we set $S(x,y)=+\infty$. This concludes our solution to the subproblem $\mathrm{T}(x,y)$. The problem $T(x,y,z)$ was constructed to simplify the presentation of our recursive solution.
@@ 4.3 Solving the Original Problem in Trees
This section demonstrates how to solve the original MSCS problem for the tree $\mathrm{T}$. Let $S$ denote an optimal solution to this problem. If $\mathrm{T}$ is monochromatic, we return any vertex as a solution. Assume $\mathrm{T}$ has more than one color. Also, assume that $L$ is a leaf block of $\mathrm{T}$, and $\mathrm{N}$ is its only neighboring block. Let $y$ be the vertex of $\mathrm{N}$ adjacent to a vertex of $L$. By Lemma 3.2, any optimal solution has exactly one vertex, say $x$, from L. Thus, $x$ is the closest vertex of $S$ to all vertices in $L$. Hence the original problem is an instance of the $\mathrm{T}(x,y)$ problem that was introduced earlier. Since we do not know $x$, we try all vertices of $L$. Therefore,